OpenStudy (anonymous):

PLEASE HELP? What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

hartnn (hartnn):

can u find the slope of line y = 4x – 2

OpenStudy (anonymous):

4?

hartnn (hartnn):

that is correct.

hartnn (hartnn):

and slope of parallel lines are ..... ?

OpenStudy (anonymous):

they're the same but have different y-intercept

hartnn (hartnn):

correct again, so you got the slope of your line as m=4 the equation you will use is y=mx+c any ideas how u find c ?

OpenStudy (anonymous):

no not rly.. i usually use the equation: y-y1=m(x-x1) to solve it.

hartnn (hartnn):

yeah, u can use that also, what u get when u use that equation ?

OpenStudy (anonymous):

y-6= 4(x - -3)

hartnn (hartnn):

u know what to substitute for x1 and y1 ?

OpenStudy (anonymous):

yeah y1=6 x1=-3

OpenStudy (anonymous):

oh nvm I got the answer! Thnaks for your help though(: